OpenStudy (anonymous):

Find an equation of a parabola that has the given numbers as solutions: f(2/3)=0 f(2)=0

OpenStudy (anonymous):

(x-2)(3x-2) = 0

OpenStudy (anonymous):

or x^2 - 8x + 4 = 0

OpenStudy (anonymous):

Awesome thanks!

OpenStudy (anonymous):

How about (2-3i)(2+3i)

OpenStudy (anonymous):

a multiple of this equation would also have same roots eg 2x^2 - 16x + 8 = 0

OpenStudy (anonymous):

Thanks, gave you a medal. :)

OpenStudy (anonymous):

thats not a parabola - expanded that comes to 4 + 9 = 13

OpenStudy (anonymous):

you welcome